OpenStudy (anonymous):

What is the meaning of the word "accessible" as it is used in the following sentence? Jarrod had no accessible proof that his little sister hid his car keys, but he couldn't think of anyone else who would have put them in the dog's bowl. Respectful Critical Available Forgetful Read the following excerpt. The politician tried to refute all the allegations against him. Unfortunately for him, though, the evidence was far too compelling. As much as he tried, he couldn't sway voter opinion, and he lost the election. Using the context, the word compelling means outrageous boring convincing thoughtful Read the following excerpt. Using the context of the sentence, select the word that best completes the sentence. We may try in __________ to solve our problems on our own, but it is not until we face the truth that we will learn that we must work together for progress. credibility predict vain forgive

OpenStudy (anonymous):

1. Available 2. Outrageous (MAYBE) Not sure of the last one. Sorry, I'm not much help...

OpenStudy (compassionate):

\(\Huge{\color{purple}{\textbf{W}} \color{orange}{\cal{E}} \color{green}{\mathbb{L}} \color{blue}{\mathsf{C}} \color{maroon}{\rm{O}} \color{red}{\tt{M}} \color{gold}{\tt{E}} \space \color{orchid}{\mathbf{T}} \color{Navy}{\mathsf{O}} \space \color{OrangeRed}{\boldsymbol{O}} \color{Olive}{\mathbf{P}} \color{Lime}{\textbf{E}} \color{DarkOrchid}{\mathsf{N}} \color{Tan}{\mathtt{S}} \color{magenta}{\mathbb{T}} \color{goldenrod}{\mathsf{U}} \color{ForestGreen}{\textbf{D}} \color{Salmon}{\mathsf{Y}} \ddot \smile }\) 1. Critical 2. Convincing 3. Vain 4.

OpenStudy (anonymous):

wooo thats kool thats too :)

OpenStudy (anonymous):

thanxs

OpenStudy (anonymous):

W/E/L/C/O/M/E /T/O /O/P/E/N/S/T/U/D/Y/

OpenStudy (compassionate):

\(\Huge\color{red}{\mathbb{LOL~FAIL!!!}}\)

OpenStudy (anonymous):

lol how do you do that ???